OpenStudy (anonymous):

24 coins. 1/3rd are pennies, 1/4th are nickels, and 1/6th are dimes. How many are quarters? and how much money in all? full step by step help please!!!!

OpenStudy (anonymous):

Well theres only one time of coin left? So you have to add up 1/3, 1/4 and 1/6 then take that away from 1. This will give you a fraction which will be the number of quarters left. Then from this just use the fractions and the knowledge you have 24 coins and voila. You have your answer

OpenStudy (anonymous):

multiply 24 to each fraction and add all those numbers together. Those are the number of coins that are not quarters then simply just subtract it from 24

OpenStudy (radar):

8 pennies 8 cents 6 nickels 30 cents 4 dimes 40 cents 6 quarters $1.50 total $2.28

OpenStudy (radar):

These figures were obtained by the "Romero" method.

OpenStudy (radar):

Using the dimc88 method. number of quarters = 1-1/3-1/4-1/6=24/24-8/24-6/24-4/24=6/24=1/4 were quarters.

OpenStudy (yuki):

YAY